OpenStudy (anonymous):

high school physics: A marshmallow is dropped from a 5-meter high pedestrian bridge and 0.83 seconds later, it lands right on the head of an unsuspecting person walking underneath. How tall is the person with the marshmallow on his head?

OpenStudy (amistre64):

meters need gravity as -4.9

OpenStudy (amistre64):

h(t) = -.5g t^2 +Vi t + Hi h(t) = -4.9 t^2 +0 t + 5 h(.83) = -4.9(.83)^2 + 5 = 1.62439; this is how far its dropped in .83 seconds

OpenStudy (amistre64):

5 - 1.62439 = 3.37561 meters to go; so id say the person is: 3.37561 meters tall

OpenStudy (anonymous):

Thanks for breaking it down for me. It has been way too long since I took physics!

OpenStudy (amistre64):

this is about the only physics i know how to do :) youre welcome